OpenStudy (chris215):

Find f '(−4), if f(x) = (5x^2 + 6x)(3x^2 + 7). Round your answer to the nearest integer. Use the hyphen symbol, -, for negative values.

OpenStudy (chris215):

i got -3214

OpenStudy (matt101):

I get a different answer. What is you general equation for f'(x)?

OpenStudy (chris215):

f'(x) = (10x+6)*(3x^2+7) + (5x^2 + 6x)*(6x) f'(-4) = (-40+6)*(48+7) + (80-24)*(-24) = -34*55 + 56*(-24)

OpenStudy (matt101):

Woops the answer you got is exactly right! I accidentally used 4 instead of -4!

OpenStudy (chris215):

thats ok thank you!